INAF LBC+MODS queue observing, 2012 January 15-16
 
Observer:  K. Boutsia, R. Speziali 
Telescope Operator: G. Bechetti 
Support Astronomer: O. Kuhn 
Instrument: LBC+MODS 
 Summary 
DOME CLOSED because of snow storm.
Since we are closed, we decided to obtain all necessary calibrations for the projects using MODS
  Data Summary  
  Details 
04:48 Starting up MODS 
-  mods1 start gui 
  -  execMODS ~MODSeng/modsScripts/Support/modsWake.pro 
  -  modsDisp 
 
 
04:50 Bias frames  
-  execMODS allbias.cal  
mods1b.20120116./mods1r.20120116 0001-0030
 
 
05:45 Nonino MOS calibrations 
-  execMODS gmosFlatNonino0647.cal - flats with mask ID520958 vflat 7 
mods1b / mods1r 0031-0035 (clear - GG495) 
 
 
05:55 Power outage... the compressor went off. we might need to restart everything 
-  we go on with the calibrations, since according to support astronomer and TO internal mods calibrations should not be affected.
  -  execMODS gmosARCNonino0647.cal - ARCS with mask ID520958 
mods1b / mods1r 0036-0038
  -  execMODS gmosARCNonino0717.cal - ARCS with mask  ID532613 
mods1b / mods1r 0039-0041
  -  execMODS gmosFlatNonino0717.cal - flats with mask ID532613 vflat 7 
mods1b / mods1r 0042-0046 (clear - GG495) 
  -  execMODS gmosFlatNonino0744.cal - flats with mask ID581623 vflat 7 
mods1b / mods1r 0047-0051 (clear - GG495) 
  -  when this script finished the red images did not appear in newdata and in the control panel it was saying - last image 48 - next image 51, the status during the execution of the OB also showed no errors. In order to retrieve those images and make them appear in /newdata we had to type in the control panel - command line box: red fitsflush  
  -  execMODS gmosARCNonino0744.cal - ARCS with mask  ID581623 
mods1b / mods1r 0052-0054 
 
 
07:16 Calibrations for standard stars longslits 
-  execMODS grlamps5.0.cal -  ARCS with LS60x5 for both channels using template script 
mod1b / mod1r 0055-0057
  -  execMODS SlitFlatLBT5.0.cal - flats with LS60x5 for both channels using template script 
mod1b / mod1r 0058-0062 (clear - GG495)
 
 
07:35 Another power glitch - we go on with mods.  
-  execMODS SlitFlatLBT1.0.cal - flats with LS5x60x1.0 for both channels using template script 
mod1b / mod1r 0063-0067 (clear - GG495)
  -  execMODS grlamps1.0.cal -  ARCS with LS5x60x1.0 for both channels using template script 
mod1b / mod1r 0068-0070
  -  execMODS  SlitLESSFlat.cal - slitless flats for pixel to pixel response - we used template script 
mod1b 0071-0080 / mod1r 0071-0075 
 
 
08:28 Fiore MOS calibrations - scripts have been slightly changed from what the PI has sent 
-  execMODS grmosflats.56.FIORE1.cal - we have changed the archive info and nimg from 1 to 5 
mod1b 0081-0090 / mod1r 0076-0085 
  -  some slits (probably the ref boxes) in the red flats for the QTH1 are saturated 
  -  execMODS grmosflats.57.FIORE2.cal - we have changed the archive info and nimg from 1 to 5 
mod1b 0091-0100 / mod1r 0086-0095 
  -  execMODS grmosflats.55.FIORE3.cal - we have changed the archive info and nimg from 1 to 5 
mod1b 0101-0110 / mod1r 0096-0105 
  -  some slits (probably the ref boxes) are saturated. This happens more often in red but ut also happens occasionally in the blue (QTH1+ UG5) 
 
 
09:46 We have modified the PIs OBs because he was using the prism template... instead we have used the grlamps.cal template after replacing the slit ID with the mosID 
-  need to ask Olga for updated templates
  -  execMODS grmoslamps.55.FIORE3.cal  
mod1b 0111-0113 / mod1r 0106-0108 
  -  execMODS grmoslamps.56.FIORE1.cal  
mod1b 0114-0116 / mod1r 0109-0111 
  -  execMODS grmoslamps.57.FIORE2.cal  
mod1b 0117-0119 / mod1r 0112-0114 
  -  execMODS grpixflats.cal 
mod1b 0120-0129 / mod1r 0115-0119 
  -  execMODS grpixflats.cal 
mod1b 0130-0139 / mod1r 0120-0124 repeat with INAF in the pi_name
  -  execMODS bias8K.cal 
mod1b 0140-0149 / mod1r 0125-0134
 
 
11:30 execMODS ~MODSeng/modsScripts/Support/modsSleep.pro  
 
-- 
KonstantinaBoutsia - 16 Jan 2012